OpenStudy (anonymous):

Can somebody check a few answers for me? Please I really need to get these right. :/

OpenStudy (anonymous):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x − 4. I said: y = 3x − 7

OpenStudy (anonymous):

@Hero @amistre64 @e.mccormick ?

OpenStudy (e.mccormick):

Well, know how to test these?

OpenStudy (anonymous):

well I actually graphed them, if that's what you mean?

OpenStudy (e.mccormick):

No, an easy way to mathematical test is to plug it in and solve to see if it is true. (3, 2) and y = 3x − 7 2 = 3(3) - 7 If you solve that and it is true, then you did it right because your equations do have the same slope, which means hey are parallel.

OpenStudy (anonymous):

i'm not quite sure I get it. I thought since they both have 3 (or 3/1) as their slop, they'd be parallel?

OpenStudy (e.mccormick):

2 = 3(3) - 7 2 = 9 - 7 2 = 2 Because 2 does equal 2, this is true. I put in the x and y values and got a true result, therefore the equation works. That, plus the slope part, means the answer is correct. What if I had (2,4) as the point, but the equation was the same: 4 = 3(3) - 7 4 = 9 - 7 4 = 2 That is false. That is what would happen if you had done it wrong. This is why a mathematical check is useful.

OpenStudy (anonymous):

ohhh. ok. that makes sense. thank you

OpenStudy (anonymous):

would you mind looking at more?

OpenStudy (e.mccormick):

Sure. If they are the same sort of thing, you can try doing the checks and if you have any issues I can help. If it is something different, I can show you how to work or check it.

OpenStudy (anonymous):

they're different, though a few are somewhat similar

OpenStudy (anonymous):

The question was: What is the length of the portion BC of the cloth? Options: 8 cos 35° sin 35 degrees by 8 cos 35 degrees by 8 8 sin 35°

OpenStudy (anonymous):

I said 8 sin 35degrees

OpenStudy (e.mccormick):

So you used the definition of sine and a little algebra, right?

OpenStudy (anonymous):

Well, i didn't really use the formal process I should have. That was the only option that seemed to fit

OpenStudy (anonymous):

^ B and C were meant to be sin 35 degrees/ 8 and cos 35 degrees/ 8

OpenStudy (e.mccormick):

Hehe. OK. Know your trig ratios? Heard of SOH CAH TOA? What that is short for is this: \(\sin = \dfrac{opp}{hyp}\) \(\cos = \dfrac{adj}{hyp}\) \(\tan = \dfrac{opp}{adj}\) If you use that, then you see that: \(\sin 35^\circ = \dfrac{opp}{8}\) If you solve for opp, or the opposite side (which is BC) then you get the answer you selected.

OpenStudy (anonymous):

ok. awesome :)

OpenStudy (e.mccormick):

If you can remember SOH CAH TOA and the meanings of opposite, adjacent, and hypotenuse you can set up any trig ratio very quickly.

OpenStudy (anonymous):

ok. I should probably have just worked that one out normally. But I don't feel all too comfortable with things outside sin/cos/tan. It wasn't explained well to me. Which is revelavant in this one> The next question was: What is the distance, in feet, that the box has to travel to move from point A to point C? Options: 12 by sec 65 degrees 12 cosec 65° (what I chose) 12 sin 65° 12 by cot 65 degrees I don't feel all to confortable with things outside sin/cos/tan

OpenStudy (e.mccormick):

OK, so of opposite, adjacent, and hypotenuse, what is side AC?